An experiment is performed aboard the International Space Station to verify that linear momentum is conserved during collisions in a zero-g environment. The experiment involves a 3-D completely inelastic collision of three drops of honey. At the moment just before they all collide, the masses and velocities of the drops are

What is your prediction for the speed, V, of the combined honey drop after the collision? An experiment is performed aboard the International Space Station to verify that linear momentum is conserved during collisions in a zero-g environment. The experiment involves a 3-D completely inelastic collision of three drops of honey. At the moment just before they all collide, the masses and velocities of the drops are m,-347 ,-139 m/s* m,-52.3 ,-(12.5 m/sj m,-76.5 g18.3 m/s v, - 13.9 m/slr What is your prediction for the speed, V, of the combined honey drop after the collision? Number m s
Answer #1

Here ,

let the speed of combined honey drop is V

m1 * v1 + m2 * v2 + m3 * v3 = (m1 + m2 + m3) * V

34.7 * 13.9 i + 52.3 * 12.5 j + 76.5 * 18.3 k = (34.7 + 52.3 + 76.5) * V

V = 2.95 i + 4 j + 8.6 k m/s

V = sqrt(2.95^2 + 4^2 + 8.6^2) = 9.9 m/s

the combined speed after the collision is 9.9 m/s